OpenStudy (anonymous):

write the equation of the line that passes through 2 2 and 6 3 in standard form

OpenStudy (anonymous):

\[\frac{y-3}{3-2}=\frac{x-6}{6-2}\]

OpenStudy (anonymous):

y-3=(x-6)/4 4y-12=x-6 x-4y+6=0